I am doing triprotic acid titrations.
I have an imaginary acid (H3A) 10mL with the concentration 0.760M and we are titrating a NaOH with the concentration of 0.320M.
I have to find the amount of NaOH in mol and mL for
1. first half-equivalence point
2.first equivalence point
3. second-half equivalence point
4.second equivalence point
5. third half-equivalence point and
6. third equivalence point.

and then I have to find the pH for each of these points.
1
Expert's answer
2020-03-10T09:07:48-0400

amount of NaOH in mol for every point: (0.76/100)/(2/№point)= 0.0038; 0.0076; 0.0114; 0.0152; 0.019; 0.0228

amount of NaOH in mL for every point: n/0.000320= 11.875; 23.75; 35.63; 47.5; 59.38; 71.25

pH for every point: pH=-lg(CH3AK1+CH2A+K2+CHAK3)1/2